HOW DO I WORK 6 3/10 + 3 2/5=
Answer by josmiceli(19441)

6 3/10 + 3 2/5
note that 2/5 = 4/10 so,
6 3/10 + 3 4/10
Now just add the fractions
3/10 + 4/10 = 7/10
Now add the whole numbers
6 + 3 = 9
Now add fraction + whole number
9 + 7/10 = 9 7/10